# TakuNet: an Energy-Efficient CNN for Real-Time Inference on Embedded UAV systems in Emergency Response Scenarios

### v1

@@ 3 TakuNet model
The architecture of the proposed TakuNet is depicted in Figure 1. Our choices are driven by the need to develop a model that combines strong performance in image classification with limited computational demands, enabling real-time execution on embedded hardware. From a general point of view, we take advantage of depth-wise convolutions and grouped point-wise convolutions, inspired by [17] to reduce the number of convolutional layer parameters. We harness the Global Response Normalization layer [50] to promote feature diversity and increase contrast and selectivity along channels, enabling the model to identify and differentiate more effectively between relevant features. In addition, dense connections [19] are used to encourage feature reuse and better gradient propagation. Finally, similarly to [24], we have included layers of dilated convolutions that favor a sudden increase in the receptive field without enlarging the number of parameters. They also allow the model to learn features on multiple scales, enabling it to capture more complex representations.
@@ 3.1 TakuNet Architecture
in which a depth-wise convolution with stride $1$ is applied. The dilation and padding are set to $1$ for all Taku Block of each stage, except for the second depthwise convolution where they are set to $2$. The four stages have a depth of $(5,5,5,4)$, respectively. The number of stages and blocks are empirically set, aiming for the best tradeoff between accuracy on one side and model size and efficiency on the other side. At the end of each stage, a Downsampler block is responsible for reducing the size of the feature maps, while expanding the number of channels. Inspired by [19] and [54], the feature maps produced by the last Taku Block are concatenated and mixed with those coming from the dense connection. To further reduce the number of parameters, we apply a grouped point-wise convolutional layer [40] on the new concatenated feature maps, where the number of groups is given by $\left\lfloor\nicefrac{{(\text{in\_channels}+\text{out\_channels})}}{{4}}\right\rfloor$. Here $in\_channels$ refers to the number of channels of the feature maps input to the stage, while $out\_channels$ refers to the number of channels of the feature maps output to the last Taku Block. Next, a Batch Normalization and a ReLU6 activation function are applied. In the first 3 stages of the architecture, the output is then dimensionally scaled by a max-pooling layer, while in the last stage average pooling is applied. Each of these pooling layers is then followed by a Global Response Normalization (GRN) [50] function.
@@ 4.2 Aerial Image Disaster Event Recognition Version 2 (AIDERv2) Dataset
Model Parameters Model Size (MB) F1-score (%) Compute FLOPs
@@
Table 3: Comparison of model parameters, model memory footprint, F1-score, and FLOPs on the latest available version of the AIDER [23] dataset. Elements highlighted in bold refer to the best results, while those underlined refer to the second-best results.
@@ 4.4 Evaluation and Performance Metrics
Model Parameters Model Size (MB) F1-score (%) Compute FLOPs
@@
Table 4: Comparison of Model Parameters, Size, F1-score, and FLOPs on AIDERv2. Elements highlighted in bold refer to the best results, while those underlined refer to the second-best results.
@@ 5 Results
Both EmergencyNet and TinyEmergencyNet are re-implemented using the PyTorch [^3] https://pytorch.org framework, adhering rigorously to the instructions detailed in the respective articles [24, 33]. During this process, we observe minor discrepancies in the publicly available implementation of EmergencyNet, while no official implementation of TinyEmergencyNet is found online. Other architectures included in the comparison are sourced directly from the TorchVision library [^4] https://pytorch.org/vision/stable/index.html, ensuring consistency and reproducibility in benchmarking. All methodologies in comparison are trained for 300 epochs with a batch size of 64 on AIDER (Table 3) and AIDERv2 (Table 4) datasets, employing the training settings specified in the respective publications.
@@
Table 5: Model performance evaluation on embedded platforms, with fps calculated as the mean latency over multiple runs with batch size equal to 1. TakuNet’s fps on Jetson Orin are reported after TensorRT optimization.
@@ 5.1 Results on AIDER
We first perform a set of experiments on AIDER [23] to validate the effectiveness of our method against some of the well-known CNNs, along with EmergencyNet and TinyEmergencyNet. Leveraging the stem to reduce the size of the input images by a factor of 4 has significantly enhanced the efficiency of our approach. Although our method employs a sparse integration of features along the channel axis, dense channel features integration enables TakuNet to perform proficiently on AIDER while considerably reducing its parameter count. The results shown in Table 3 highlight that TakuNet is capable of providing the same F1-score as EmergencyNet while using the minimum number of parameters and FLOPs among the various competitors.
These results point out that our method is the best in terms of the trade-off between F1-score/FLOPs, making a significant contribution to high-efficiency aerial image classification methods. Furthermore, when TakuNet is trained using floating point 32 precision, it can achieve the best F1-score among competitors, making it the new state-of-the-art on AIDER. This result is important for two other reasons. First, the difference in F1-score between the use of 16-bit floating-point precision (FP=16) and 32-bit floating-point precision (FP=32) is minimal. However, the efficiency gap between the two is substantial, primarily due to hardware acceleration, enabling the optimization of half-precision computation. Secondly, not all processors support hardware acceleration for 16-bit floating-point precision. In such cases, the version trained with 32-bit floating-point precision can be employed as an alternative.
@@ 5.2 Results on AIDERv2
Deeper experiments on AIDERV2 showcase the effectiveness of a few architectural design choices. In fact, by removing dense connections the model does not converge during training, while by omitting GRN the overall F1-score decreases to $0.942$. The Dilated convolutions and the refiner block also are key elements for the high-quality outputs of TakuNet  since their exclusion causes the accuracy to drop at $0.953$ and $0.950$, respectively.
@@ 6 Conclusion
Our proposed TakuNet achieves significant results, providing high accuracy compared to competing approaches while maintaining a minimal number of parameters, FLOPs, and memory usage. This makes it a versatile solution capable of addressing a wide range of requirements in embedded applications, particularly in emergency response scenarios. Additionally, we emphasize the significant impact of hardware-specific characteristics on model performance, such as the varying latencies of embedded devices when processing specific model components. By focusing on hardware-aware design, we demonstrate that not only can impressive performance levels be achieved thanks to hardware acceleration, but also that the accuracy gap between FP16 and FP32 precision is negligible.

### v3

@@ 3 TakuNet model
The architecture of the proposed TakuNet is depicted in Figure 1. Our choices are driven by the need to develop a model that combines strong performance in image classification with limited computational demands – i.e. a limited number of parameters and floating point operations – enabling real-time execution on embedded hardware.
From a general point of view, we take advantage of several elements introduced in the recent literature. The first is the depth-wise convolutions and grouped point-wise convolutions, inspired by [17, 40] to reduce the number of convolutional layer parameters. We harness the Global Response Normalization layer [50] to promote feature diversity and increase contrast and selectivity along channels, enabling the model to identify and differentiate more effectively between relevant features. In addition, dense connections [19] are used to encourage feature reuse and better gradient propagation. Finally, following  [24], we have included a layer of dilated convolution in the stem to capture more distant features without enlarging the number of parameters.
@@ 3.1 TakuNet Architecture
in which a depth-wise convolution with stride $1$ is applied. The dilation and padding are set to $1$ for all Taku Block of each stage. The four stages have a depth of $(5,5,5,4)$, respectively. The number of stages and blocks are empirically set, aiming for the best tradeoff between accuracy on one side and model size and efficiency on the other side. At the end of each stage, a Downsampler block is responsible for reducing the size of the feature maps, while expanding the number of channels. Inspired by [19, 54], the feature maps produced by the last Taku Block are concatenated and mixed with those coming from the dense connection. To further reduce the number of parameters, we apply a grouped point-wise convolutional layer [40] on the new concatenated feature maps, where the number of groups is given by $\left\lfloor\nicefrac{{(\text{in\_channels}+\text{out\_channels})}}{{4}}\right\rfloor$. Here $in\_channels$ refers to the number of channels of the feature maps input to the stage, while $out\_channels$ refers to the number of channels of the feature maps output to the last Taku Block.
Next, a Batch Normalization and a ReLU6 activation function are applied. In the first 3 stages of the architecture, the output is then dimensionally scaled by a max-pooling layer, while in the last stage average pooling is applied. Each of these pooling layers is then followed by a Global Response Normalization (GRN) [50] function.
@@ 4.2 Aerial Image Disaster Event Recognition Version 2 (AIDERv2) Dataset
Model Parameters Model Size (MB) F1-score FLOPS
@@
Table 3: Comparison of model parameters, memory footprint, F1-score, and FLOPs on the latest AIDER [23] version, where images are set to $240\times 240$ pixels. Elements highlighted in bold refer to the best results, while those underlined refer to the second-best results.
@@ 4.4 Evaluation and Performance Metrics
Model Parameters Model Size (MB) F1-score FLOPS
@@
Table 4: Comparison of Model Parameters, Size, F1-score, and FLOPs on AIDERv2 [42], where image height and width are set to $224$. Elements highlighted in bold refer to the best results, while those underlined refer to the second-best results.
@@ 5 Results
Both EmergencyNet and TinyEmergencyNet are re-implemented using the PyTorch [^3] https://pytorch.org framework, adhering rigorously to the instructions detailed in the respective articles [24, 33]. During this process, we observe minor discrepancies in the publicly available implementation of EmergencyNet, while no official implementation of TinyEmergencyNet is found online. Other architectures included in the comparison are sourced directly from the TorchVision library [^4] https://pytorch.org/vision/stable, ensuring consistency and reproducibility in benchmarking. All methodologies in comparison are trained for 300 epochs with a batch size of 64 on AIDER (Table 3) and AIDERv2 (Table 4) datasets, employing the training settings specified in the respective publications.
@@
Table 5: Model performance evaluation on embedded platforms, with fps calculated as the mean latency over multiple runs with batch size equal to 1 and image size of $240$ pixels. TakuNet’s fps on Jetson Orin are reported after TensorRT optimization.
@@ 5.1 Results on AIDER
We first perform a set of experiments on AIDER [23] to validate the effectiveness of our method against some of the well-known CNNs, along with EmergencyNet and TinyEmergencyNet. Leveraging the stem to reduce the size of the input images by a factor of 4 has significantly enhanced the efficiency of our approach. Although our method employs a sparse integration of features along the channel axis, dense channel features integration enables TakuNet to perform proficiently on AIDER while considerably reducing its parameter count. The results shown in Table 3 highlight that TakuNet is capable of surpassing EmergencyNet in F1-score while using the minimum number of parameters and FLOPs among the various competitors.
These results point out that our method is the best in terms of the trade-off between F1-score/FLOPs, making a significant contribution to high-efficiency aerial image classification methods. Furthermore, when TakuNet is trained using floating point 32 precision, the F1-score obtained on the test set is lower to a minor extent. This result is important for two other reasons. First, the difference in F1-score between the use of 16-bit floating-point precision (FP=16) and 32-bit floating-point precision (FP=32) is moderated. However, the efficiency gap between the two is substantial, primarily due to hardware acceleration, enabling the optimization of half-precision computation. Secondly, not all processors support hardware acceleration for 16-bit floating-point precision. In such cases, either lossless 32-bit floating-point weight casting or 16-bit floating-point emulation can be employed.
@@ 5.2 Results on AIDERv2
Deeper experiments on AIDERV2 showcase the effectiveness of a few architectural design choices. In fact, by removing dense connections the model does not converge during training, while by omitting GRN the overall F1-score decreases to $0.950$. The channel shuffle and the refiner block also are key elements for the high-quality outputs of TakuNet  since their exclusion causes the accuracy to drop at $0.956$ and $0.955$, respectively.
@@ 6 Conclusion
Our proposed TakuNet achieves significant results, providing high accuracy compared to competing approaches while minimizing the parameter count, FLOPs, and memory usage. This makes it a versatile solution capable of addressing a wide range of requirements in embedded applications, particularly in emergency response scenarios. Additionally, we emphasize the significant impact of hardware-specific characteristics on model performance, such as the varying latencies of embedded devices when processing specific model components. By focusing on hardware-aware design, we demonstrate that impressive performance can be achieved through hardware acceleration, and that models trained with FP16 precision can be highly competitive. Such an approach is advisable for edge device models.